Fresh Water, Inc. sold an issue of 3-year $1,000 par value bonds to the public. The bonds have a 9.39 percent coupon rate and pay interest annually. The current market rate of interest on the Fresh water, Inc. bonds is 8.55 percent. What is the current market price of the bonds?
Answer #1

Annual coupon=1000*9.39%=$93.9

Hence current value=Annual coupon*Present value of annuity factor(8.55%,3)+$1000*Present value of discounting factor(8.55%,3)

=93.9*2.551734108+$1000*0.781826733

=$1021.43(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=93.9[1-(1.0855)^-3]/0.0855

=93.9*2.551734108

2.Present value of discounting factor=1000/1.0855^3

=1000*0.781826733